OpenStudy (anonymous):

Write the equation of a line that passes through the point (1,-3) and is parallel to the line 7x + y = -5. A. y = -7x + 4 B. y = -4x + 7 C. y = 4x + -7 D. y = 28x

OpenStudy (anonymous):

a

OpenStudy (anonymous):

parrallel is the same slope

OpenStudy (anonymous):

Write the equation of a line that passes through the point (-2, 1) and has a slope of 4. A. y = 4x − 7 B. y = 4x + 1 C. y = 4x + 3 D. y = 4x + 9

OpenStudy (anonymous):

plug in xy for those values to find b

OpenStudy (anonymous):

so 1=4(-2)+b

OpenStudy (anonymous):

so d is answer